### v1

# Large Language Models as Neurolinguistic Subjects: Identifying Internal Representations for Form and Meaning

Abstract
This study investigates the linguistic understanding of Large Language Models (LLMs) regarding signifier (form) and signified (meaning) by distinguishing two LLM assessment paradigms: psycholinguistic and neurolinguistic. Traditional psycholinguistic evaluations often reflect statistical biases that may misrepresent LLMs’ true linguistic capabilities. We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pair and diagnostic probing to analyze activation patterns across model layers. This method allows for a detailed examination of how LLMs represent form and meaning, and whether these representations are consistent across languages. Our contributions are three-fold: (1) We compare neurolinguistic and psycholinguistic methods, revealing distinct patterns in LLM assessment; (2) We demonstrate that LLMs exhibit higher competence in form compared to meaning, with the latter largely correlated to the former; (3) We present new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable reasoning, linguistic, arithmetic, and other cognitive abilities. The advent of LLMs has reignited cross-disciplinary discussions about what sorts of behavior are “intelligence”, even if the intelligence exhibited by LLMs may differ from human intelligence Sejnowski (2023). LLMs have drawn the attention of researchers from various fields, including linguistics, cognitive science, computer science and neuroscience, who investigate how LLMs develop and exhibit these capabilities.
Figure 1: Illustration of LLMs prcocessing the same signified (meaning) across different signifiers (forms).
Figure 2: Psycholinguistic vs. Neurolinguistic Paradigm. Both direct probability measurement and metalinguistic prompting can be considered as psycholinguistic methods, while minimal pair probing He et al. (2024) and other diagnostic probing are neurolinguistic.
There is currently a heated debate about whether LLMs understand human language or whether their performance is simply the product of complex statistical relationships Mitchell and Krakauer (2023). Some assert that LLMs have demonstrated the potential for artificial general intelligence Bubeck et al. (2023). According to Manning (2022), optimism grows that emerging knowledge-imbued systems have exhibited a degree of general intelligence. Others remain skeptical Bender and Koller (2020). Browning and LeCun (2022) argue that LLMs are limited to a superficial understanding. Bender et al. (2021) suggests that LLMs cannot possess true understanding because they lack mental models of the world and experiential knowledge.
A central aspect of this debate concerns the nature of LLMs’ linguistic representations. Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how do form and meaning intertwist with each other. Is LLM’s understanding of language meaning merely a statistical outcome based on their grasp of language form? When different languages express a shared concept with distinct forms, do LLMs create similar representations for these variations? How can we better understand the representations of form and meaning in these systems that support the observed patterns of performance?
The underlying processes remain unclear due to the opaque nature of neural networks. Therefore, we need appropriate methods to assess their true linguistic understanding.
Drawing inspirations from cognitive study on human language processing, we propose that assessment of LLMs can be divided into two primary paradigms: psycholinguistic and neurolinguistic. As illustrated in Figure 2, psycholinguistic paradigm measures the model’s output probabilities, directly reflecting the model’s behavior and performance. Neurolinguistic paradigm delves into the internal representations of LLMs.
While prior studies have treated language models as psycholinguistic subjects Futrell et al. (2019), this research is the first to propose viewing LLMs as neurolinguistic subjects as well.
When treating LLMs as psycholinguistic subjects, their responses often lean towards signifiers (form) due to their reliance on statistical correlations. This enables LLMs to produce structurally coherent language but not necessarily semantically accurate responses, as their “understanding” is shaped by statistical patterns rather than true conceptual processing Harnad (1990); Bender and Koller (2020); Nie et al. (2024), which makes psycholinguistic evaluation results might not accurately reflect the true linguistic understanding of LLMs.
In contrast, examining LLMs as neurolinguistic subjects focuses on internal representations, which may transcend surface-level statistical biases. To achieve neurolinguistic assessing, we adapted the decoding probing method presented by He et al. (2024), which we refer to as “minimal pair probing” to distinguish it from conventional diagnostic probing Belinkov and Glass (2019); Belinkov (2022). Using this method, we can examine how LLMs represent both form and meaning in a fine-grained, layer-by-layer manner.
In order to address questions about whether LLMs maintain consistent underlying representations of the same concept when form changes across multiple languages, we also create a multilingual minimal pair dataset (COMPS-ZH for Chinese and COMPS-DE for German).
By evaluating LLMs in both psycholinguistic and neurolinguistic paradigm, our results reveal:1) Psycholinguistic and neurolinguistic results reveal very different patterns, suggesting both paradigms are necessary for a comprehensive understanding of LLMs. 2) LLMs demonstrate easier, earlier and better competence in language form than meaning. 3) As linguistic form varies across different languages, the LLMs’ understanding of meaning appears to shift accordingly with the meaning competence being linearly correlated to form. Results together suggest that signifier and signified in LLMs might not be independent to each other. Maintaining a conceptual representation might rely on statistical correlations based on linguistic form.
2 Psycholinguistic vs. Neurolinguistic Paradigm
2.1 Cognitive Science Background
Psycholinguistics and neurolinguistics offer distinct yet complementary perspectives on human language processing. Psycholinguistics focuses on the psychological and cognitive processes that enable humans to understand and use language Field (2004); Traxler and Gernsbacher (2011). In contrast, neurolinguistics explores the underlying neural mechanisms and brain structures involved in language processing Friederici (2011); Brennan (2022); Kemmerer (2022). Both paradigms offer a valuable model for probing the linguistic capacities and potential intelligence of LLMs.
2.2 In LLM Assessment Research
Psycholinguistic paradigm: direct probability measurement and metalinguistic prompting
Recent studies often use prompting to evaluate the linguistic capabilities of LLMs. These implicit tests were referred to as metalinguistic judgments by Hu and Levy (2023). However, it is important to note that the performance of LLMs in specific linguistic prompting tasks only indirectly reflects their internal linguistic representations due to the inherent limitations of such prompting tasks: an LLM chat system might give a “reasonable” response just because of the statistical relationships between prompt and reply Hofstadter (1995). Hu and Levy (2023) argue that it is uncertain whether the LLMs’ responses to metalinguistic prompting align with the underlying internal representations.
Computing a model’s probability of generating two minimally different sentences is one way to address these concerns Hu and Levy (2023). The minimal difference between the two sentences (e.g. replacement of a single word) makes one sentence acceptable while the other is not Linzen et al. (2016). Here are two examples for testing grammatical and conceptual understanding, respectively:
(1) Simple agreement Warstadt et al. (2020):
       a. The cats annoy Tim. (acceptable)
       b. *The cats annoys Tim. (unacceptable)
(2) Concept understanding Misra et al. (2023):
       a. A whisk adds air to a mixture. (acceptable)
       b. *A cup adds air to a mixture. (unacceptable)
A language model is considered to perform correctly on this task if it assigns a higher probability to the acceptable sentence compared to the unacceptable one Marvin and Linzen (2018). Researchers created syntactic, semantic/conceptual, and discourse inference tasks for the minimal pair method. They provide more precise insights into the abilities of LLMs compared to metalinguistic prompting Futrell et al. (2019); Gauthier et al. (2020); Hu et al. (2020); Warstadt et al. (2020); Beyer et al. (2021); Misra et al. (2023); Kauf et al. (2023).
Through either metalinguistic judgement or direct probability measurement methods, these tasks essentially treat LLMs as psycholinguistic subjects Futrell et al. (2019). This research paradigm resembles cognitive psychology by having LLMs perform tasks, such as cloze and question answering, and then evaluating their performance without examining the internal representations. Information about the inner workings of a model is inferred either from its output or from the probabilities it assigns to different possible outputs. The internal states of the LLM (i.e. its intermediate layers) are not examined.
Neurolinguistic paradigm: diagnostic probing
Another line of research focuses on studying the internal representations, emphasizing a neurolinguistic approach to understanding LLMs. Essentially, previous diagnostic probing methods in evaluating language models can be considered as neurolinguistic paradigms as they examine the internal states of LMs Belinkov and Glass (2019); Belinkov (2022), while the term ‘neurolinguistic’ hasn’t been applied to the field before. Diagnostic probing involves training a classifier to predict linguistic properties from the hidden states of LMs. Following this paradigm, researchers decode syntactic, semantic, morphological, and other linguistic properties from the hidden states of LMs Köhn (2015); Gupta et al. (2015); Shi et al. (2016); Tenney et al. (2019); Manning et al. (2020).
3 Minimal Pair Probing = Minimal Pair + Diagnostic Probing
Minimal Pair Duality Language # of Pair Description
BLiMP Form English 67, 000 67 tasks across 12 grammatical phenomena
CLiMP Form Chinese 16, 000 16 tasks across 9 grammatical phenomena
DistilLingEval Form German 8, 000 8 German grammatical phenomena
COMPS Meaning English 49, 340 4 types of conceptual relationship
COMPS-ZH Meaning Chinese 49, 340 4 types of conceptual relationship
COMPS-DE Meaning German 49, 340 4 types of conceptual relationship
Table 1: Overview of datasets in our study.
While prior neurolinguistic approaches have explored internal representations, they often employed coarse-grained datasets and primarily focused on decoding linguistic labels from embeddings, providing a general perspective on the linguistic features encoded in LMs. In contrast, the minimal pair probing method presented by He et al. (2024) integrates minimal pair design with diagnostic probing. This combination leverages the granularity of minimal pair design and the layer-wise insights of diagnostic probing, thereby enabling a more detailed analysis of internal patterns for form and meaning. We adopt the minimal pair decoding method as the neurolinguistic paradigm in our work.
Specifically, given an LLM $f:x_{0,1,...,i}\rightarrow x_{i+1}$ trained on dataset $\mathcal{D}_{O}$, we can extract the hidden state representations $f_{l}(S)$ of the $l$-th layer of stimuli $S$. Given a minimal pair dataset $\mathcal{D}_{m}$ = $\{(S_{+}^{i},S_{-}^{i}),(z_{+}^{i},z_{-}^{i})\}$ with each sentence $S$ has a label $z$, we have internal representation $f_{l}(S_{+}^{i})$ and $f_{l}(S_{-}^{i})$ for each sentence. A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$. Note that what we focus on is the performance of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, instead of the capability of the supervised classifier $g$. To eliminate the influence of $g$, we assign a random vector $r$ to each sentence to replace the hidden states $f_{l}(S)$, and calculate the performance of the classifier $g$ on the random vector $r$ as $\mathrm{Perf}(g,\mathcal{D}_{m})$. This “random” baseline evaluates the predictability of the probing classifier itself. Finally, the performance of the LLM $f$ is defined as follows:
$\mathrm{Perf}(f,\mathcal{D}_{O})\triangleq\frac{\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})-\mathrm{Perf}(g,\mathcal{D}_{m})}{1-\mathrm{Perf}(g,\mathcal{D}_{m})}$ (1)
4 Experiment Setup
4.1 Datasets
We use minimal pair probing for English, Chinese, and German to assess grammaticality (form) and conceptuality (meaning). Table 1 presents the overall dataset information used in our experiments. Further dataset descriptions are in Appendix B.
4.2 Models
In our experiments, we used three open-source LLMs, two English-centric LLMs (Llama2 and Llama3), and one multilingual LLMs (Qwen) with a focus on English and Chinese. These models were trained on different amounts of English, Chinese, and German data (see Table 2).
Resource Level Llama2 Llama3 Qwen
English High High High
Chinese Mid Mid High
German Low Low Low
Table 2: Resource level for different languages across three LLMs. Note the resource levels are qualitative assessments based on available information, as specific quantitative data is not provided by the developers.
Llama2 and Llama3
Llama2 (Touvron et al., 2023b) and Llama3 (AI, 2024) are two English-centric LLMs which represent an advanced iteration of the Llama foundation models developed by Meta AI (Touvron et al., 2023a). The Llama madels were trained on publicly available corpora predominantly in English. Despite this focus, Llama models are also exposed to a limited amount of multilingual data. Llama 1, for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages. Llama 2 expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data. Therefore, English-centric models harness multilingual abilities (Lai et al., 2023). In this work, we use Llama2-7B and Llama3-8B for our experiments.
QWen
QWen is a series of LLMs developed by Alibaba Inc. Bai et al. (2023). Qwen was trained on 2-3 trillion tokens of multilingual pretaining data. It is essentially a multilingual LLM with the focus on English and Chinese. We use the Qwen-7B model in our experiments.
Duality Method Example
Form Direct {Mice are hurting a waiter, Mice was hurting a waiter}
Meta Here are two English sentences: 1) Mice are hurting a waiter. 2) Mice was hurting a waiter. Which
sentence is a better English sentence? Respond with either 1 or 2 as your answer. Answer: {1, 2}
Meaning Direct {Helmet can absorb shocks, Cap can absorb shocks}
Meta What word is most likely to come next in the following sentence (helmet, or cap)? What can absorb
shocks? {helmet, cap}
Table 3: Prompt examples for baseline methods. Region where we measure probability is marked in color. Correct sentences and answers are in blue; incorrect in red.
4.3 Setup for Psycholinguistic Paradigms
Direct
Direct probability measurement calculates the probability of entire sentences based on model logits. Accuracy is determined by whether the model assigns a higher probability to the grammatically or conceptually correct sentence within the minimal pair.
Meta
Metalinguistic prompting involves explicitly asking a question or specifying a task that requires a judgement about a linguistic expression. Following Hu and Levy (2023), we use a single prompt each minimal pair to present both sentences at once. For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences, and compare the probabilities assigned by the model to each answer option, “1” or “2”. For meaning tasks, we reformulate the property into a question and compare the probabilities of acceptable and unacceptable concepts as sentence continuations. Table 3 presents the prompts used in the experiments.
4.4 Setup for Minimal Pair Probing
Sentence Embedding
We extract the last token in each sentence from each layer to serve as the representation for the whole sentence. Last token pooling ensures the representation contains the information of all preceding tokens Meng et al. (2024).
Probing Performance
We use logistic regression as the probing classifier and F1 score as the evaluation metrics. The score for $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$ and $\mathrm{Perf}(g,\mathcal{D}_{m})$ is calculated as the average F1 score across 5 cross-validation folds. Final performance $\mathrm{Perf}(f,\mathcal{D}_{O})$ is given by Formula 1.
Saturation and Maximum Layer
We define the feature learning Saturation Layer as the layer where performance first reaches 95% of the peak on the curve. This layer indicates the number of layers required for the model to adequately learn specific linguistic features, after which its ability to capture these features stabilizes. The Maximum Layer is the layer at which performance reaches its peak.
5 Results
5.1 Psycholinguistic vs Neurolinguistic
Figure 3: Psycholinguistic vs. Neurolinguistic Results. Direct refers to direct probability measurement; meta refers to metalinguistic prompting and neuro refers to minimal pair probing, which is a neurolinguistic paradigm.
In order to compare psycholinguistic and neurolinguistic paradigms, we use the last layer’s performance in the neuro method  [^1] We refer to minimal pair probing as the neuro method for simplicity in the results section., as both direct probability measurement and metalinguistic prompting rely on the last layer of LLMs.
Differences among Direct, Meta, and Neuro results
Figure 3 shows distinct results for Direct, Meta, and Neuro, with a general performance order of Neuro > Direct > Meta. This suggests that relying on a single method may not give a comprehensive assessment of language models, and it is best to employ multiple methods for evaluation.
Neuro is similar to Direct on meaning but not on form.
The Neuro method consistently outperforms others across models when assessing form, often nearing 1.0, while Direct and Meta show more variability and generally lower performance. The Neuro method performs similarly to Direct in meaning evaluation, yet remains significantly higher than Meta.
Meta method performs consistently poorly.
This serves as a caution for the current trend of instruction-based prompting, suggesting that Meta’s limitations could reflect issues with instruction prompting’s effectiveness.
Instruction tuning influences LLMs’ Meta performance on form but not much on meaning.
Chat version models’ meta performance on form varies a lot compared to their base versions, but meaning understanding keeps stable.
5.2 Detailed Neurolinguistic Results within English
LLMs incrementally encode grammar and concepts.
Figure 4: Detailed English minimal pair probing results. The pink, yellow, and blue curves represent three grammatical levels: morphology, the semantics-syntax interface, and syntax, respectively. The red curves denote conceptual tasks.
To investigate how LLMs encode form and meaning in a fine-grained and layer-by-layer manner, we split form back into 12 original grammatical tasks and meaning into 4 original conceptual tasks. Figure 4 indicates that Llama-2 rapidly learns grammatical and conceptual information in the first layers incrementally. Performance improves across layers and stabilizes at some point. Similar results for all 6 models can be found in Figure 14 of Appendix D.
Figure 4 shows result curves for 16 specific linguistic tasks. 12 curves (with pink, yellow, and blue colors) visualize the results of grammatical tasks, while the 4 red curves represent the results of the conceptual tasks.
LLMs encode grammatical features better than conceptual features.
Figure 5: English minimal pair probing results for form and meaning.
The results in Figure 5 show that the performance scores for conceptual understanding (red curves) are significantly lower than those for grammatical understanding (blue curves). This pattern is consistent across all six models, suggesting a universal characteristic of LLMs.
LLMs encode meaning after form.
Figure 6: Feature learning saturation (95% of peak) and maximum layer.
We compute the feature learning saturation and maximum layers for all 12 grammatical tasks and 4 conceptual tasks, averaging them to represent form and meaning, respectively. From Figure 6, we can see that the saturation and maximum layers for meaning are generally higher than those for form across all six models. This suggests that LLMs stabilize their encoding of grammatical features before conceptual features, indicating that LLMs learn the form of language first and then the meaning of language.
Instruction tuning won’t change internal linguistic representation much.
Figure 7: Performance difference between the base and the instruction versions of LLMs. The shaded areas indicate standard deviation where $\text{std}_{\text{diff}}=\sqrt{\text{std}_{\text{base}}^{2}+\text{std}_{\text{chat}}^{2}}$.
From Figure 7, instruction tuning does not significantly alter the internal linguistic representations of the models. The performance differences for both form and meaning remain close to zero throughout all layers, suggesting that the overall linguistic encoding remains stable despite the tuning.
5.3 Meaning Encoding across Different Forms
How does LLMs’ understanding of meaning change when the form (language) varies? Since our COMPS-ZH and COMPS-DE datasets align with the concepts in the English COMPS dataset, we can explore whether LLMs’ grasp of different linguistic forms for the same concept correlates with their understanding of meaning across languages.
Figure 8: Minimal pair probing results of all three languages.
Our previous results and the detailed results in Figures 15 and 16 suggest that instruction tuning has little influence on the internal representations. Therefore we focus on the base LLMs in the following analysis.
Cross-lingual Pattern
From Figure 8, for all models and languages, form consistently achieves higher F1 scores than meaning, indicating it’s easier for LLMs to make stronger grasp of structural elements compared to conceptual comprehension.
Form needs less data to capture compared to meaning.
From Table 2, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
Figure 9: Meaning competence of LLMs is correlated to their form competence.
Meaning competence is correlated to form competence.
Figure 9 illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs.The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
6 Discussion
Form and Meaning: observations from Saussure’s semiotics Through various experiments focused on Saussure’s concepts of the signifier (form) and the signified (meaning), we have identified several key phenomena. First, LLMs consistently learn linguistic form before they grasp meaning. This suggests a developmental trajectory where statistical patterns in syntax and grammar are more readily captured by the model than conceptual understanding. Second, the models’ formal competence is generally superior to their semantic competence. This is evident in their ability to decode grammaticality structures accurately but with less reliable conceptual accuracy.
Moreover, we observe a linear correlation between form and meaning competence, especially when linguistic forms vary across languages while the underlying meaning remains constant. This pattern implies that an LLM’s understanding of meaning is heavily dependent on its grasp of form. In cases where the form changes (such as between languages), meaning appears to follow the variations in form, indicating that conceptual representation in LLMs is anchored to formal structures rather than an independent comprehension of meaning. These findings highlight the models’ reliance on statistical correlations for understanding meaning, further emphasizing their dependence on formal patterns rather than direct, intrinsic understanding of natural language semantics.
These results help explain the phenomenon of LLMs generating “confidently incorrect” responses, commonly referred to as hallucination, has been a significant issue for a long time Ji et al. (2023). Our results provide a theoretical explanation for this phenomenon from a semiotic and neurolinguistic perspective. Our findings also align with previous studies Mahowald et al. (2024), which suggests that LLMs have good formal competence, but their performance on functional competence (understanding language in the world) is spotty.
Developmental difference between human and machine intelligence. From a perspective of developmental psychology, human kids typically acquire conceptual understanding before mastering grammar Bloom (2002); Tomasello (2005). Pinker (2009)’s semantic bootstrapping hypothesis posits that children initially learn vocabulary through semantic information and then use this semantic knowledge to infer syntactic structures. In contrast, our results indicate that LLMs learn grammar before meaning. Human intelligence is a combination of statistical inference and causal reasoning, whereas LLMs’ intelligence is more likely a result of statistical inference Tenenbaum et al. (2011); Gopnik and Wellman (2012); Lake et al. (2017). Given this nature, the fact that LLMs learn form first might be because grammatical patterns are easier to statistically capture compared to meaning.
Symbol grounding problem and the quest for human-like intelligence
In human language, the relationship between the signifier and the signified is often flexible and context-dependent, allowing for a more independent connection between syntax and semanticsHarnad (1990). Human cognitive development typically involves acquiring conceptual understanding first, followed by the learning of rules and syntax. In contrast, our study shows that LLMs grasp syntax before meaning, relying on statistical correlations within formal structures to infer semantic content. This difference highlights a fundamental divergence between human and machine intelligence, as LLMs do not possess an inherent understanding of meaning detached from the formal structures they analyze.
These observations suggest that, for LLMs to develop human-like intelligence, they must transcend mere statistical pattern recognition. This will likely require the integration of world knowledge and grounded experiences that go beyond linguistic inputs. To achieve a more robust form of artificial intelligence that mirrors human cognition, models must be able to ground symbols in real-world contexts, establishing a basis for genuine understanding Tenenbaum et al. (2011); Lake et al. (2017) As it stands, the symbol grounding problem remains a significant barrier, and addressing it may be essential for constructing systems capable of true human-like reasoning and understanding.
7 Conclusion
By analyzing Large Language Models (LLMs) as both psycholinguistic and neurolinguistic subjects, we find that LLMs excel at encoding linguistic form but show limitations in understanding meaning. Their grasp of meaning is largely dependent on statistical associations with form. By introducing a cognitive neuroscience perspective, along with semiotics, we hope will inspire further research to deepen our understanding of the language capabilities of LLMs.
Limitations
We didn’t include experiments covering more languages, which might limit the generalizability of our findings.
The evaluation results for German are poor, which might be attributed to the issue of very long sentences in the DistilLingEval dataset.
Due to limitations in computational resources, we did not perform our experiments on larger-size LLMs, which may lead our results to be biased towards small-scale models. Further studies should include larger models to validate and potentially generalize our findings.
Ethics Statement
Our project may raise awareness within the computational linguistics community about those low-resource languages.
Appendix A More Psycholinguistic and Neurolinguistic Paradigms’s Cognitive Science Background
Psycholinguistics often involves examining real-time language processing, linguistic knowledge storage, and language acquisition, using behavioral experimental methods such as reading times and eye-tracking. Neurolinguistics, on the other hand, focuses on the neural basis of language, employing techniques such as functional magnetic resonance imaging (fMRI), electroencephalography (EEG), and magnetoencephalography (MEG) to map linguistic functions to specific brain regions and to investigate how neural activity correlates with linguistic tasks.
While psycholinguistics aims to reveal mental processes underlying language use, neurolinguistics seeks to uncover the neural pathways that implement these processes.
Our minimal pair probing is inspired by cognitive neuroscience. In the field of neurolinguistics, decoding analysis has become a fundamental technique in cognitive neuroscience. It tries to extract information encoded in neural patterns (Kriegeskorte et al., 2006). It trains a classifier to predict properties of the stimulus (e.g. a particular image or word input), from the neural responses. If the accuracy of the trained classifier is significantly better than chance, we conclude that the neural data encodes information about the predicted stimulus properties (Norman et al., 2006; Haynes and Rees, 2006).
Mesgarani and Chang (2012) is a representative work employing decoding analysis in the realm of language and speech. They use electrocorticography (ECoG) to record neural responses from subjects who are listening to speech. Leveraging decoding analysis, they were able to differentiate between neural patterns induced by attended speech and those elicited by background speech, (to be ignored by the test persons), thereby highlighting the perceptual differences between the two speech stimuli conditions.
While psycholinguistic approaches provide valuable insights into LLMs’ functional capabilities, they often fall short in revealing the underlying mechanisms of language processing. The opaque nature of neural network structures means that performance on external tasks does not necessarily reflect the internal cognitive processes at play. This gap necessitates a neurolinguistic approach to gain a deeper understanding of how LLMs encode and process language.
Appendix B Dataset Details
For each language, we use one dataset for grammatical minimal pairs and one dataset for conceptual minimal pairs.
B.1 Form: BLiMP, CLiMP, and DistilLingEval
BLiMP
BLiMP (Warstadt et al., 2020) is a comprehensive English dataset of grammatical minimal pairs. It consists of minimal pairs for 13 higher-level linguistic phenomena in the English language, further divided into 67 distinct realizations, called paradigms. Each paradigm comprises 1,000 individual minimal pairs, resulting in a total corpus size of 67,000 data points.
CLiMP
CLiMP (Xiang et al., 2021) is a corpus of Chinese grammatical minimal pairs consisting of 16 datasets, each containing 1,000 sentence pairs. CLiMP covers 9 major Chinese language phenomena in total, less than the BLiMP dataset due to the less inflectional nature of Mandarin Chinese. The vocabulary of the CLiMP dataset is based on the translation of the BLiMP dataset, with words and features specific to Chinese added.
DistilLingEval
DistilLingEval (Vamvas and Sennrich, 2021) is a dataset of German grammatical minimal pairs. It consists of minimal pairs for eight German linguistic phenomena. This dataset contains 82,711 data samples in total.
B.2 Meaning: COMPS, COMPS-ZH, and COMPS-DE
COMPS
COMPS (Misra et al., 2023) is an English dataset of conceptual minimal pairs for testing an LLM’s knowledge of everyday concepts (e.g., a beaver/*gorilla has a flat tail). This dataset contains 49,340 sentence pairs, constructed using 521 concepts and 3,592 properties. Concepts in the pairs constitute 4 types or knowledge relationship: taxonomy, property norms, co-occurrence and random.
COMPS-ZH and COMPS-DE
COMPS-DE and COMPS-ZH are newly developed datasets featuring conceptual minimal pairs in Chinese and German, derived from the English COMPS dataset (Misra et al., 2023). In the realm of multilingual NLP research, it is a common practice to extend English datasets to other languages using human translation, machine translation, or translation assisted by LLMs (Nie et al., 2023; Wang et al., 2023; Beniwal et al., 2024).
In this study, to create COMPS-DE and COMPS-ZH from the original English COMPS, we employed a hybrid approach that integrated process machine translation with meticulous human verification.
Specifically, we translated the concepts and properties of the English COMPS individually, subsequently merging them to form complete sentences and compose conceptual minimal pairs. The translation process began with the use of the Google Translate API [^2] https://cloud.google.com/translate, which provided initial translations of concepts and properties into German and Chinese.
Following this, native speakers of Chinese and German manually checked and refined these translations to ensure accuracy and quality. The manual review emphasized two main areas: accuracy of concepts and grammatical consistency of properties. For concepts, the focus was on correcting ambiguities that might arise from machine translation. For properties, attention was given to maintaining grammatical consistency with the original English text, such as ensuring subject-verb agreement, which is particularly challenging in German translations.
In summary, out of 521 concepts, manual corrections were made to 57 entries in the Chinese dataset and 49 in the German dataset. Similarly, from 3,592 properties, 713 required manual corrections in the Chinese dataset, and 512 in the German dataset. This rigorous process was essential for preserving the integrity and reliability of the translated datasets.
Appendix C More Detailed Minimal Pair Probing Results for English
Figure 10: English minimal pair probing results.
We averaged the results of 4 tasks from the same linguistic level, respectively, and show the result curves in Figure 10. Some noteworthy points from this figure include:
1) The average performances on morphological tasks (purple curve) and on tasks from the semantics-syntax interface (yellow curve) are lower than those for syntactic tasks (blue curve), suggesting that more abstract linguistic levels are easier for LLMs to learn, which also aligns with He et al. (2024). For the ellipsis task especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
2) Among the conceptual tasks, the random relationship shows significantly higher accuracy compared to the other three conceptual relationships, suggesting that LLMs find it challenging to distinguish between similar concepts.
Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much, but will learn concepts better and faster.
Figure 11: Performance difference between Llama3 and Llama2.
Figure 12: Layer-wise t-test results. T-tests were first performed separately on each linguistic task, and then Stouffer’s Z-score method Stouffer et al. (1949) was employed to aggregate the final p-value at the condition level.
Figure 11 shows the layer-wise performance differences between Llama3 and Llama2, as well as between their chat versions. The red curves (meaning) exhibit a notable positive difference in the early layers, indicating that Llama3 has better conceptual learning capabilities compared to Llama2. The blue curves (form) remain close to zero across all layers, suggesting that there is no significant improvement in grammatical capabilities in Llama3 compared to Llama2. The t-test statistics in Figure 12 further support these results.
Figure 13: T-test results between Llama2 and Llama3 feature learning saturation layer. The symbols ‘***’, and ‘*’ denote t-test p-values less than 0.001, and 0.05, respectively.
Figure 13 compares the feature learning saturation layers between Llama2 and Llama3. The results for form learning (blue bars) do not differ significantly between Llama2 and Llama3 and weakly significantly between Llama2_chat and Llama3_chat. However, the results for meaning learning (blue bars) are both highly significant, indicating that Llama3 requires fewer layers to encode conceptual features than Llama2. This suggests that Llama3 comprehends sentences faster.
Appendix D Multilingual Analysis
1) Llama’s performance on Chinese. Despite Chinese not being the primary training language of the Llama2 models, they still perform well in encoding Chinese form/grammar. However, both Qwen, which is primarily trained on Chinese, and the Llama models show relatively poor performance in understanding Chinese meaning/concepts.
2) Improvement in Llama3 for Chinese Semantics. Llama3 shows some improvement over Llama2 in understanding Chinese semantics, as indicated by the slightly higher red curve in the middle row. The improvement in syntactic understanding is minimal.
3) Qwen’s Faster Syntax Learning but Slower Semantic Encoding for Chinese. Compared to the Llama models, Qwen learns Chinese grammar faster, as indicated by the sharper rise of the blue curve. However, it encodes Chinese semantics more slowly, evidenced by the larger gap between the form and meaning curves in the early layers.
4) Poor Performance for German. For German, a low-resource language, all three models perform poorly. Despite Chinese not being a primary training language for the Llama models, their performance is relatively decent, suggesting that the actual proportion of German training data might be much smaller. This highlights differences in the resource allocation for the three languages.
Figure 14: Detailed English decoding results on 6 models.
Figure 15: Detailed Chinese decoding results on 6 models. Notice that the pink, orange, blue curves don’t denote morphology or semantics as those in English do. They are made just to make it easier to distinguish in the figure. All non-red curves represent grammatical tasks and red curves represent conceputal tasks.
Figure 16: Detailed German decoding results on 6 models. All non-red curves are grammatical tasks, and red curves are conceptual tasks.

### v2

@@
-Large Language Models as Neurolinguistic Subjects: Identifying Internal Representations for Form and Meaning
+Large Language Models as Neurolinguistic Subjects: Discrepancy in Performance and Competence for Form and Meaning
 Abstract
 This study investigates the linguistic understanding of Large Language Models (LLMs) regarding signifier (form) and signified (meaning) by distinguishing two LLM assessment paradigms: psycholinguistic and neurolinguistic.
-Traditional psycholinguistic evaluations often reflect statistical biases that may misrepresent LLMs’ true linguistic capabilities.
+Traditional psycholinguistic evaluations often reflect statistical rules that may not accurately represent LLMs’ true linguistic competence.
 We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pair and diagnostic probing to analyze activation patterns across model layers.
 This method allows for a detailed examination of how LLMs represent form and meaning, and whether these representations are consistent across languages.
-Our contributions are three-fold: (1) We compare neurolinguistic and psycholinguistic methods, revealing distinct patterns in LLM assessment; (2) We demonstrate that LLMs exhibit higher competence in form compared to meaning, with the latter largely correlated to the former; (3) We present new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets.
+We found: (1) Psycholinguistic and neurolinguistic methods reveal that language performance and competence are distinct; (2) Direct probability measurement may not accurately assess linguistic competence; (3) Instruction tuning won’t change much competence but improve performance; (4) LLMs exhibit higher competence and performance in form compared to meaning.
+Additionally, we introduce new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets.
 1 Introduction
@@
 The advent of LLMs has reignited cross-disciplinary discussions about what sorts of behavior are “intelligence”, even if the intelligence exhibited by LLMs may differ from human intelligence Sejnowski (2023).
-LLMs have drawn the attention of researchers from various fields, including linguistics, cognitive science, computer science and neuroscience, who investigate how LLMs develop and exhibit these capabilities.
-Figure 1: Illustration of LLMs prcocessing the same signified (meaning) across different signifiers (forms).
-Figure 2: Psycholinguistic vs.
+LLMs have drawn the attention of researchers from various fields, including linguistics, cognitive science, computer science, and neuroscience, who investigate how LLMs develop and exhibit these capabilities.
+Figure 1: Left: Research Question.
+Right: Psycholinguistic vs.
 Neurolinguistic Paradigm.
-Both direct probability measurement and metalinguistic prompting can be considered as psycholinguistic methods, while minimal pair probing He et al.
+Both direct probability measurement and metalinguistic prompting can be considered psycholinguistic methods, while minimal pair probing He et al.
 (2024) and other diagnostic probing are neurolinguistic.
 There is currently a heated debate about whether LLMs understand human language or whether their performance is simply the product of complex statistical relationships Mitchell and Krakauer (2023).
-Some assert that LLMs have demonstrated the potential for artificial general intelligence Bubeck et al.
-(2023).
-According to Manning (2022), optimism grows that emerging knowledge-imbued systems have exhibited a degree of general intelligence.
-Others remain skeptical Bender and Koller (2020).
-Browning and LeCun (2022) argue that LLMs are limited to a superficial understanding.
-Bender et al.
-(2021) suggests that LLMs cannot possess true understanding because they lack mental models of the world and experiential knowledge.
 A central aspect of this debate concerns the nature of LLMs’ linguistic representations.
-Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how do form and meaning intertwist with each other.
+Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how form and meaning intertwist with each other.
 Is LLM’s understanding of language meaning merely a statistical outcome based on their grasp of language form?
@@
 The underlying processes remain unclear due to the opaque nature of neural networks. Therefore, we need appropriate methods to assess their true linguistic understanding.
-Drawing inspirations from cognitive study on human language processing, we propose that assessment of LLMs can be divided into two primary paradigms: psycholinguistic and neurolinguistic.
-As illustrated in Figure 2, psycholinguistic paradigm measures the model’s output probabilities, directly reflecting the model’s behavior and performance.
-Neurolinguistic paradigm delves into the internal representations of LLMs.
-While prior studies have treated language models as psycholinguistic subjects Futrell et al. (2019), this research is the first to propose viewing LLMs as neurolinguistic subjects as well.
-When treating LLMs as psycholinguistic subjects, their responses often lean towards signifiers (form) due to their reliance on statistical correlations.
-This enables LLMs to produce structurally coherent language but not necessarily semantically accurate responses, as their “understanding” is shaped by statistical patterns rather than true conceptual processing Harnad (1990); Bender and Koller (2020); Nie et al.
-(2024), which makes psycholinguistic evaluation results might not accurately reflect the true linguistic understanding of LLMs.
-In contrast, examining LLMs as neurolinguistic subjects focuses on internal representations, which may transcend surface-level statistical biases.
-To achieve neurolinguistic assessing, we adapted the decoding probing method presented by He et al.
-(2024), which we refer to as “minimal pair probing” to distinguish it from conventional diagnostic probing Belinkov and Glass (2019); Belinkov (2022).
-Using this method, we can examine how LLMs represent both form and meaning in a fine-grained, layer-by-layer manner.
-In order to address questions about whether LLMs maintain consistent underlying representations of the same concept when form changes across multiple languages, we also create a multilingual minimal pair dataset (COMPS-ZH for Chinese and COMPS-DE for German).
-By evaluating LLMs in both psycholinguistic and neurolinguistic paradigm, our results reveal:1) Psycholinguistic and neurolinguistic results reveal very different patterns, suggesting both paradigms are necessary for a comprehensive understanding of LLMs.
-2) LLMs demonstrate easier, earlier and better competence in language form than meaning.
-3) As linguistic form varies across different languages, the LLMs’ understanding of meaning appears to shift accordingly with the meaning competence being linearly correlated to form.
-Results together suggest that signifier and signified in LLMs might not be independent to each other.
-Maintaining a conceptual representation might rely on statistical correlations based on linguistic form.
+Drawing inspirations from the cognitive study on human language processing, we propose that the assessment of LLMs can be divided into two primary paradigms: psycholinguistic and neurolinguistic.
+As illustrated in Figure 1, the psycholinguistic paradigm measures the model’s output probabilities, directly reflecting the model’s behavior and performance.
+The neurolinguistic paradigm delves into the internal representations of LLMs.
+When treating LLMs as psycholinguistic subjects, their responses may leverage their grasp of form, relying on statistical correlations, to create an illusion of understanding meaning.
+This enables LLMs to produce structurally coherent but not necessarily semantically accurate responses, as their “understanding” is shaped by patterns rather than true conceptual processing Harnad (1990); Bender and Koller (2020); Nie et al.
+(2024).
+Consequently, psycholinguistic evaluations tend to reflect performance rather than competence, as they assess external outputs that may not fully capture the underlying linguistic knowledge encoded within the model.
+This mismatch suggests that psycholinguistic evaluation results might not accurately represent the true linguistic competence of LLMs.
+In contrast, examining LLMs as neurolinguistic subjects focuses on internal representations, providing a more direct assessment of competence by moving beyond surface-level biases Firestone (2020).
+To achieve this, we adapted the decoding probing method by He et al.
+(2024), referred to as “minimal pair probing”, to analyze how LLMs encode form and meaning across layers.
+This approach allows for a finer distinction between performance and competence, revealing insights that psycholinguistic methods might overlook.
+In order to address questions about whether LLMs maintain consistent underlying representations of the same concept when the form changes across multiple languages, we also create a multilingual minimal pair dataset (COMPS-ZH for Chinese and COMPS-DE for German).
+By evaluating LLMs in both psycholinguistic and neurolinguistic paradigms, we found: 1) Psycholinguistic and neurolinguistic results reveal very different patterns, suggesting both paradigms are necessary for a comprehensive understanding of LLMs.
+2) Though more intrinsic than metalinguistic prompting, direct probability measurement may still not accurately assess linguistic competence, as it remains influenced by statistical patterns.
+3) LLMs acquire competence in linguistic form more easily, earlier, and with greater accuracy than in meaning.
+4) As linguistic form varies across languages, LLMs’ understanding of the same concept shifts accordingly, with meaning competence linearly correlated to form.
+This suggests that signifier and signified in LLMs may not be independent, and maintaining conceptual representations likely depends on statistical correlations with form.
 2 Psycholinguistic vs. Neurolinguistic Paradigm
@@
 A language model is considered to perform correctly on this task if it assigns a higher probability to the acceptable sentence compared to the unacceptable one Marvin and Linzen (2018).
-Researchers created syntactic, semantic/conceptual, and discourse inference tasks for the minimal pair method.
+Researchers have created syntactic, semantic/conceptual, and discourse inference tasks for the minimal pair method.
 They provide more precise insights into the abilities of LLMs compared to metalinguistic prompting Futrell et al.
@@
 (2019).
-This research paradigm resembles cognitive psychology by having LLMs perform tasks, such as cloze and question answering, and then evaluating their performance without examining the internal representations.
+This research paradigm resembles cognitive psychology by having LLMs perform tasks, such as cloze and question answering, and then evaluating their performance without examining the internal representations, in a manner similar to how subjects participate in psychological experiments.
 Information about the inner workings of a model is inferred either from its output or from the probabilities it assigns to different possible outputs.
@@
 Another line of research focuses on studying the internal representations, emphasizing a neurolinguistic approach to understanding LLMs.
-Essentially, previous diagnostic probing methods in evaluating language models can be considered as neurolinguistic paradigms as they examine the internal states of LMs Belinkov and Glass (2019); Belinkov (2022), while the term ‘neurolinguistic’ hasn’t been applied to the field before.
+Essentially, diagnostic probing methods in evaluating language models can be considered as neurolinguistic paradigms as they examine the internal states of LMs Belinkov and Glass (2019); Belinkov (2022), while the term ‘neurolinguistic’ hasn’t been applied to the field before.
 Diagnostic probing involves training a classifier to predict linguistic properties from the hidden states of LMs.
@@
 Table 1: Overview of datasets in our study.
-While prior neurolinguistic approaches have explored internal representations, they often employed coarse-grained datasets and primarily focused on decoding linguistic labels from embeddings, providing a general perspective on the linguistic features encoded in LMs.
-In contrast, the minimal pair probing method presented by He et al.
-(2024) integrates minimal pair design with diagnostic probing.
-This combination leverages the granularity of minimal pair design and the layer-wise insights of diagnostic probing, thereby enabling a more detailed analysis of internal patterns for form and meaning.
-We adopt the minimal pair decoding method as the neurolinguistic paradigm in our work.
-Specifically, given an LLM $f:x_{0,1,...,i}\rightarrow x_{i+1}$ trained on dataset $\mathcal{D}_{O}$, we can extract the hidden state representations $f_{l}(S)$ of the $l$-th layer of stimuli $S$.
-Given a minimal pair dataset $\mathcal{D}_{m}$ = $\{(S_{+}^{i},S_{-}^{i}),(z_{+}^{i},z_{-}^{i})\}$ with each sentence $S$ has a label $z$, we have internal representation $f_{l}(S_{+}^{i})$ and $f_{l}(S_{-}^{i})$ for each sentence.
-A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$.
-Note that what we focus on is the performance of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, instead of the capability of the supervised classifier $g$.
-To eliminate the influence of $g$, we assign a random vector $r$ to each sentence to replace the hidden states $f_{l}(S)$, and calculate the performance of the classifier $g$ on the random vector $r$ as $\mathrm{Perf}(g,\mathcal{D}_{m})$.
-This “random” baseline evaluates the predictability of the probing classifier itself.
-Finally, the performance of the LLM $f$ is defined as follows:
-$\mathrm{Perf}(f,\mathcal{D}_{O})\triangleq\frac{\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})-\mathrm{Perf}(g,\mathcal{D}_{m})}{1-\mathrm{Perf}(g,\mathcal{D}_{m})}$ (1)
-4 Experiment Setup
-4.1 Datasets
-We use minimal pair probing for English, Chinese, and German to assess grammaticality (form) and conceptuality (meaning).
-Table 1 presents the overall dataset information used in our experiments.
-Further dataset descriptions are in Appendix B.
-4.2 Models
-In our experiments, we used three open-source LLMs, two English-centric LLMs (Llama2 and Llama3), and one multilingual LLMs (Qwen) with a focus on English and Chinese.
-These models were trained on different amounts of English, Chinese, and German data (see Table 2).
-Resource Level Llama2 Llama3 Qwen
-English High High High
-Chinese Mid Mid High
-German Low Low Low
-Table 2: Resource level for different languages across three LLMs.
-Note the resource levels are qualitative assessments based on available information, as specific quantitative data is not provided by the developers.
-Llama2 and Llama3
-Llama2 (Touvron et al., 2023b) and Llama3 (AI, 2024) are two English-centric LLMs which represent an advanced iteration of the Llama foundation models developed by Meta AI (Touvron et al., 2023a).
-The Llama madels were trained on publicly available corpora predominantly in English.
-Despite this focus, Llama models are also exposed to a limited amount of multilingual data.
-Llama 1, for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages.
-Llama 2 expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data.
-Therefore, English-centric models harness multilingual abilities (Lai et al., 2023).
-In this work, we use Llama2-7B and Llama3-8B for our experiments.
-QWen
-QWen is a series of LLMs developed by Alibaba Inc.
-Bai et al.
-(2023).
-Qwen was trained on 2-3 trillion tokens of multilingual pretaining data.
-It is essentially a multilingual LLM with the focus on English and Chinese.
-We use the Qwen-7B model in our experiments.
 Duality Method Example
@@
 shocks? {helmet, cap}
-Table 3: Prompt examples for baseline methods. Region where we measure probability is marked in color. Correct sentences and answers are in blue; incorrect in red.
-4.3 Setup for Psycholinguistic Paradigms
+Table 2: Prompt examples for baseline methods. The region where we measure probability is marked in color. Correct sentences and answers are in blue; incorrect in red.
+While prior neurolinguistic approaches have explored internal representations, they often employed coarse-grained datasets and primarily focused on decoding linguistic labels from embeddings, providing a general perspective on the linguistic features encoded in LMs.
+In contrast, the minimal pair probing method presented by He et al.
+(2024) integrates minimal pair design with diagnostic probing.
+This combination leverages the granularity of minimal pair design and the layer-wise insights of diagnostic probing, thereby enabling a more detailed analysis of internal patterns for form and meaning.
+We adopt minimal pair decoding as the neurolinguistic paradigm in our work.
+Specifically, given an LLM $f:x_{0,1,...,i}\rightarrow x_{i+1}$ trained on dataset $\mathcal{D}_{O}$, we can extract the hidden state representations $f_{l}(S)$ of the $l$-th layer of stimuli $S$.
+Given a minimal pair dataset $\mathcal{D}_{m}$ = $\{(S_{+}^{i},S_{-}^{i}),(z_{+}^{i},z_{-}^{i})\}$ with each sentence $S$ has a label $z$, we have internal representation $f_{l}(S_{+}^{i})$ and $f_{l}(S_{-}^{i})$ for each sentence.
+A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$.
+Note that our focus is the performance of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, instead of the capability of the supervised classifier $g$.
+To eliminate the influence of $g$, we set up a “random” baseline, where we assign a random vector $r$ to each sentence to replace the hidden states $f_{l}(S)$ and calculate the performance of the classifier $g$ on the random vector $r$ as $\mathrm{Perf}(g,\mathcal{D}_{m})$, aiming to evaluate the inherent predictability of the probing classifier.
+Finally, the performance of the LLM $f$ is defined as follows:
+$\mathrm{Perf}(f,\mathcal{D}_{O})\triangleq\frac{\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})-\mathrm{Perf}(g,\mathcal{D}_{m})}{1-\mathrm{Perf}(g,\mathcal{D}_{m})}$ (1)
+4 Experiment Setup
+4.1 Datasets and Models
+We use minimal pair probing for English, Chinese, and German to assess grammaticality (form) and conceptuality (meaning).
+Table 1 presents the overall dataset information used in our experiments.
+We use Llama2-7B, Llama3-8B, and Qwen-7B models in both base and chat versions.
+Further dataset and model descriptions are in Appendix B and C.
+4.2 Setup for Psycholinguistic Analysis
 Direct
-Direct probability measurement calculates the probability of entire sentences based on model logits.
+Direct probability measurement calculates the probability of a sentence based on model logits.
 Accuracy is determined by whether the model assigns a higher probability to the grammatically or conceptually correct sentence within the minimal pair.
@@
 Metalinguistic prompting involves explicitly asking a question or specifying a task that requires a judgement about a linguistic expression.
-Following Hu and Levy (2023), we use a single prompt each minimal pair to present both sentences at once.
-For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences, and compare the probabilities assigned by the model to each answer option, “1” or “2”.
+Following Hu and Levy (2023), we use one prompt for a minimal pair to present both sentences at once.
+For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences and compare the probabilities assigned by the model to each answer option, “1” or “2”.
 For meaning tasks, we reformulate the property into a question and compare the probabilities of acceptable and unacceptable concepts as sentence continuations.
 Table 2 presents the prompts used in the experiments.
-4.4 Setup for Minimal Pair Probing
+Figure 2: Psycholinguistic (meta and direct) and neurolinguistic performance across models and linguistic tasks.
+The x-axis represents different models and conditions (base and chat), while the y-axis categorizes linguistic tasks based on structural (syntax, morphology, syntax-semantics interface) and conceptual (meaning) levels.
+4.3 Setup for Neurolinguistic Analysis
 Sentence Embedding
@@
 The Maximum Layer is the layer at which performance reaches its peak.
+Unsupervised Analysis
+We use t-SNE to visualize the sentence embedding of Llama2-7B for English form tasks. We employ PCA to reduce the dimensionality of the sentence embedding to 50 before applying t-SNE.
 5 Results
+Figure 3: (a) Averaged psycholinguistic (meta and direct) and neurolinguistic results across models and tasks.
+t-tests were conducted on the original (pre-averaging) results between base and chat models, with p-values annotated.
+(b) Psycholinguistic and neurolinguistic performance for form (morphology, syntax-semantics interface, and syntax) and meaning (concept).
+(c) Competence and performance gap drops after instruction tuning.
 5.1 Psycholinguistic vs Neurolinguistic
-Figure 3: Psycholinguistic vs.
-Neurolinguistic Results.
-Direct refers to direct probability measurement; meta refers to metalinguistic prompting and neuro refers to minimal pair probing, which is a neurolinguistic paradigm.
-In order to compare psycholinguistic and neurolinguistic paradigms, we use the last layer’s performance in the neuro method  [^1] We refer to minimal pair probing as the neuro method for simplicity in the results section., as both direct probability measurement and metalinguistic prompting rely on the last layer of LLMs.
-Differences among Direct, Meta, and Neuro results
-Figure 3 shows distinct results for Direct, Meta, and Neuro, with a general performance order of Neuro > Direct > Meta.
-This suggests that relying on a single method may not give a comprehensive assessment of language models, and it is best to employ multiple methods for evaluation.
-Neuro is similar to Direct on meaning but not on form.
-The Neuro method consistently outperforms others across models when assessing form, often nearing 1.0, while Direct and Meta show more variability and generally lower performance.
-The Neuro method performs similarly to Direct in meaning evaluation, yet remains significantly higher than Meta.
-Meta method performs consistently poorly.
-This serves as a caution for the current trend of instruction-based prompting, suggesting that Meta’s limitations could reflect issues with instruction prompting’s effectiveness.
-Instruction tuning influences LLMs’ Meta performance on form but not much on meaning.
-Chat version models’ meta performance on form varies a lot compared to their base versions, but meaning understanding keeps stable.
-5.2 Detailed Neurolinguistic Results within English
-LLMs incrementally encode grammar and concepts.
-Figure 4: Detailed English minimal pair probing results.
-The pink, yellow, and blue curves represent three grammatical levels: morphology, the semantics-syntax interface, and syntax, respectively.
-The red curves denote conceptual tasks.
-To investigate how LLMs encode form and meaning in a fine-grained and layer-by-layer manner, we split form back into 12 original grammatical tasks and meaning into 4 original conceptual tasks.
-Figure 4 indicates that Llama-2 rapidly learns grammatical and conceptual information in the first layers incrementally.
-Performance improves across layers and stabilizes at some point.
-Similar results for all 6 models can be found in Figure 14 of Appendix D.
-Figure 4 shows result curves for 16 specific linguistic tasks.
-12 curves (with pink, yellow, and blue colors) visualize the results of grammatical tasks, while the 4 red curves represent the results of the conceptual tasks.
-LLMs encode grammatical features better than conceptual features.
-Figure 5: English minimal pair probing results for form and meaning.
-The results in Figure 5 show that the performance scores for conceptual understanding (red curves) are significantly lower than those for grammatical understanding (blue curves).
+Figure 2 shows the performance of LLMs across all linguistic tasks.
+Figure 3-(a) demonstrates the averaged performance of LLMs across models and 4 levels (syntax, morphology, syntax-semantics interfaces, concept).
+Figure 3-(b) presents the average performance of LLMs across form and meaning tasks for Direct, Meta, and Neuro [^1] We refer to minimal pair probing as Neuro for simplicity.
+methods.
+We use the last layer’s performance in the Neuro method when comparing psycho- and neurolinguistic paradigms, as both direct probability measurement and metalinguistic prompting rely on the last layer of LLMs.
+Language performance and competence are distinct (Competence $>$ Performance).
+Figure 3 shows distinct results between language performance and competence.
+Moving from Meta → Direct → Neuro, the evaluation focus gradually shifts from language performance (task execution ability) to language competence (the underlying linguistic ability).
+Within the same task category, Neuro methods consistently yield higher performance than Direct methods, which in turn outperform Meta methods.
+This indicates that when evaluating pure linguistic competence, LLMs perform well, but their performance drops when assessed in a task-based setting.
+Tasks that emphasize language performance become more difficult, even if their language competence is high.
+For example, in the Neuro setting, performance on Syntax tasks reaches 97%, while in the Meta setting, it drops to 56.1%, showing a significant gap.
+This suggests that even when an LLM has strong competence in a given task, its performance can significantly decline when assessed under performance-oriented evaluation.
+Direct probability measurement might not be true competence assessment.
+As the Neuro method measures the internal representations of LLMs directly, it could serve as a reliable ground truth for estimating linguistic competence.
+Direct probability measurement falls short of achieving this ground truth in form assessment (especially for syntax and syntax-semantics-interface as shown in Figure 3-(b)).
+LLMs exhibit stronger mastery of form than meaning, regardless of performance or competence.
+As shown in Figure 3-(b), LLMs consistently perform better on form-related tasks than on meaning-related tasks.
+This trend holds regardless of whether the model is a base or chat version.
+Crucially, this pattern is evident across all evaluation methods.
+This indicates that LLMs have a stronger grasp of linguistic form than conceptual meaning, whether assessed through task execution or underlying capability.
+Instruction tuning won’t change much competence but improve performance.
+Neuro results between the base and chat versions of LLMs reveal that instruction fine-tuning does not significantly alter the language competence of the models (t-test between Neuro-Base vs.
+Neuro-Chat as shown in Figure 3-(a)).
+With instruction fine-tuning (chat versions of LLMs), Meta performance on form improves significantly while meaning understanding remains stable.
+Figure 3-(c) illustrates that after instruction tuning, the competence-performance gap (Neuro-Meta) significantly decreases for form-related tasks, while the change for meaning-related tasks remains relatively small.
+This indicates that fine-tuning with well-designed instructions helps LLMs improve their performance on form-related tasks, bringing them closer to their underlying competence.
+However, for meaning-related tasks, instruction tuning does not lead to a fundamental improvement in understanding.
+This indicates that more optimized information access strategies can enhance the external performance of language models, particularly for form-related tasks.
+5.2 Neurolinguistic Analysis [^2] Raw results for English, Chinese, and German can be found in Figure 8, 9 and 10 in the Appendix.
+Figure 4: t-SNE visualization of embedding differences between acceptable and unacceptable sentences, with red for syntax, purple for morphology, and yellow for the syntax-semantics interface.
+Layer-wise unsupervised dynamics reveal gradual emergence of form features
+Figure 4 illustrates the layer-wise differences between embeddings for grammatically correct and incorrect sentences.
+In early layers, the embedding difference appears scattered and unstructured, but as depth increases, they form clearer clusters, indicating a progressively refined sensitivity to syntactic correctness.
+By Layer 16 and beyond, distinct clusters emerge corresponding to syntax, morphology, and syntax-semantics interface.
+The results demonstrate that LLMs encode grammaticality judgments dynamically across layers, progressively structuring linguistic representations.
+Moreover, the formation of distinct clusters for different linguistic phenomena in the unsupervised analysis provides supporting evidence for subsequent supervised classification.
+Gradual decline in encoding performance from structure to meaning.
+The results in Figure 5-(c) show that the performance scores for conceptual understanding are significantly lower than those for grammatical understanding.
 This pattern is consistent across all six models, suggesting a universal characteristic of LLMs.
-LLMs encode meaning after form.
-Figure 6: Feature learning saturation (95% of peak) and maximum layer.
+Moreover, as illustrated in Figure 5-(a),(b), the encoding performance progressively declines from more structural tasks to more semantic tasks—spanning syntax, morphology, the syntax-semantic interface, and finally conceptual understanding.
+This highlights that LLMs encode features less effectively as the tasks shift from structure-focused to meaning-focused.
+Figure 5: (a) Neuro probing results for 12 tasks in Llama-2 (4 syntax, 4 morphology, 4 syn.-sem.
+inter., 4 concept).
+(b) Averaged performance across four linguistic levels in Llama-2.
+(c) Averaged performance on form vs.
+meaning across all six models.
+(d) Feature learning saturation (95% of peak) and maximum layer.
+(e) Perf.
+difference between base and instruction-tuned LLMs.
+(f) Neuro probing results for Chinese and German.
+(g) Correlation between meaning competence and form competence.
+LLMs encode form earlier than meaning.
 We compute the feature learning saturation and maximum layers for all 12 grammatical tasks and 4 conceptual tasks, averaging them to represent form and meaning, respectively.
-From Figure 6, we can see that the saturation and maximum layers for meaning are generally higher than those for form across all six models.
-This suggests that LLMs stabilize their encoding of grammatical features before conceptual features, indicating that LLMs learn the form of language first and then the meaning of language.
-Instruction tuning won’t change internal linguistic representation much.
-Figure 7: Performance difference between the base and the instruction versions of LLMs.
-The shaded areas indicate standard deviation where $\text{std}_{\text{diff}}=\sqrt{\text{std}_{\text{base}}^{2}+\text{std}_{\text{chat}}^{2}}$.
-From Figure 7, instruction tuning does not significantly alter the internal linguistic representations of the models.
-The performance differences for both form and meaning remain close to zero throughout all layers, suggesting that the overall linguistic encoding remains stable despite the tuning.
-5.3 Meaning Encoding across Different Forms
+As shown in Figure 5-(d), the saturation and maximum layers for meaning are generally higher than those for form across all six models.
+This suggests that LLMs stabilize their encoding of grammatical features before conceptual features.
+Instruction tuning has minimal impact on the internal linguistic representations.
+As Figure 5-(e) shows, performance differences for form and meaning remain near zero across all layers, indicating that instruction tuning minimally impacts internal linguistic representations, consistent with our psycholinguistic vs.
+neurolinguistic analysis.
+5.3 Multilingual analysis
 How does LLMs’ understanding of meaning change when the form (language) varies?
 Since our COMPS-ZH and COMPS-DE datasets align with the concepts in the English COMPS dataset, we can explore whether LLMs’ grasp of different linguistic forms for the same concept correlates with their understanding of meaning across languages.
-Figure 8: Minimal pair probing results of all three languages.
-Our previous results and the detailed results in Figures 15 and 16 suggest that instruction tuning has little influence on the internal representations.
-Therefore we focus on the base LLMs in the following analysis.
+Our previous results suggest that instruction tuning has little influence on the internal representations.
+Therefore we focus on the base LLMs here.
 Cross-lingual Pattern
-From Figure 8, for all models and languages, form consistently achieves higher F1 scores than meaning, indicating it’s easier for LLMs to make stronger grasp of structural elements compared to conceptual comprehension.
-Form needs less data to capture compared to meaning.
-From Table 2, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
-Figure 9: Meaning competence of LLMs is correlated to their form competence.
+From Figure 5-(f), for all models and languages, form consistently achieves higher performance than meaning, indicating it’s easier for LLMs to make a stronger grasp of structural elements compared to conceptual comprehension.
+Extended multilingual results can be found in Appendix E.
 Meaning competence is correlated to form competence.
-Figure 9 illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs.The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
+Figure 5-(g) illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs.
+The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
 6 Discussion
-Form and Meaning: observations from Saussure’s semiotics Through various experiments focused on Saussure’s concepts of the signifier (form) and the signified (meaning), we have identified several key phenomena.
-First, LLMs consistently learn linguistic form before they grasp meaning.
-This suggests a developmental trajectory where statistical patterns in syntax and grammar are more readily captured by the model than conceptual understanding.
+Language performance vs. competence: using neurolinguistic probing to reveal deeper competencies.
+Our results demonstrate distinct patterns between the neurolinguistic and psycholinguistic evaluation methods.
+We observed that LLM performance evaluated using the Meta method was significantly lower than that of Direct, while Direct also performed notably worse than the Neuro method.
+This finding extends Hu and Levy (2023), who compared direct probability measurement (Direct) and meta-linguistic prompting (Meta) and argued that performance does not always reflect competence.
+While Hu and Levy (2023) suggested that Direct probability measurement is a more reliable approach for assessing competence, our findings reveal that there is still a significant gap between Direct and Neuro probing, which directly examines internal representations.
+Hence, we recommend Neuro probing for evaluating competence.
+Neuro probing further validates the discrepancy between performance and competence.
+While Meta results underperform, this does not necessarily indicate that the LLMs lack the underlying linguistic competence.
+Instead, it may reflect limitations in information access, as suboptimal prompts can prevent models from exhibiting their full capabilities.
+Specifically, prompting failures do not always equate to a lack of encoded knowledge.
+This observation is consistent with previous studies Firestone (2020); Lampinen (2024) that emphasize the importance of carefully controlling performance conditions for both humans and machines.
+methods can often reveal deeper competencies within the models.
+Form and meaning: observations from Saussure’s semiotics
+Our results reveal that LLMs consistently learn linguistic form before they grasp meaning.
+This may suggest a developmental trajectory where statistical patterns in syntax and grammar are more readily captured by the model than conceptual understanding.
 Second, the models’ formal competence is generally superior to their semantic competence.
 This is evident in their ability to decode grammaticality structures accurately but with less reliable conceptual accuracy.
-Moreover, we observe a linear correlation between form and meaning competence, especially when linguistic forms vary across languages while the underlying meaning remains constant.
-This pattern implies that an LLM’s understanding of meaning is heavily dependent on its grasp of form.
-In cases where the form changes (such as between languages), meaning appears to follow the variations in form, indicating that conceptual representation in LLMs is anchored to formal structures rather than an independent comprehension of meaning.
-These findings highlight the models’ reliance on statistical correlations for understanding meaning, further emphasizing their dependence on formal patterns rather than direct, intrinsic understanding of natural language semantics.
-These results help explain the phenomenon of LLMs generating “confidently incorrect” responses, commonly referred to as hallucination, has been a significant issue for a long time Ji et al.
+We further observe a linear correlation between form and meaning competence, particularly when linguistic forms vary across languages while meaning remains constant.
+This suggests that LLMs’ understanding of meaning might rely heavily on form, with conceptual representation anchored to formal structures rather than independent meaning comprehension.
+These results offer a semiotic and neurolinguistic explanation for LLMs’ long-standing issue of generating “confidently incorrect” responses, i.e.
+hallucinations Ji et al.
 (2023).
-Our results provide a theoretical explanation for this phenomenon from a semiotic and neurolinguistic perspective.
-Our findings also align with previous studies Mahowald et al.
-(2024), which suggests that LLMs have good formal competence, but their performance on functional competence (understanding language in the world) is spotty.
-Developmental difference between human and machine intelligence.
-From a perspective of developmental psychology, human kids typically acquire conceptual understanding before mastering grammar Bloom (2002); Tomasello (2005).
-Pinker (2009)’s semantic bootstrapping hypothesis posits that children initially learn vocabulary through semantic information and then use this semantic knowledge to infer syntactic structures.
-In contrast, our results indicate that LLMs learn grammar before meaning.
-Human intelligence is a combination of statistical inference and causal reasoning, whereas LLMs’ intelligence is more likely a result of statistical inference Tenenbaum et al.
-(2011); Gopnik and Wellman (2012); Lake et al.
-(2017).
-Given this nature, the fact that LLMs learn form first might be because grammatical patterns are easier to statistically capture compared to meaning.
-Symbol grounding problem and the quest for human-like intelligence
-In human language, the relationship between the signifier and the signified is often flexible and context-dependent, allowing for a more independent connection between syntax and semanticsHarnad (1990).
-Human cognitive development typically involves acquiring conceptual understanding first, followed by the learning of rules and syntax.
-In contrast, our study shows that LLMs grasp syntax before meaning, relying on statistical correlations within formal structures to infer semantic content.
-This difference highlights a fundamental divergence between human and machine intelligence, as LLMs do not possess an inherent understanding of meaning detached from the formal structures they analyze.
-These observations suggest that, for LLMs to develop human-like intelligence, they must transcend mere statistical pattern recognition.
-This will likely require the integration of world knowledge and grounded experiences that go beyond linguistic inputs.
-To achieve a more robust form of artificial intelligence that mirrors human cognition, models must be able to ground symbols in real-world contexts, establishing a basis for genuine understanding Tenenbaum et al.
-(2011); Lake et al.
-(2017) As it stands, the symbol grounding problem remains a significant barrier, and addressing it may be essential for constructing systems capable of true human-like reasoning and understanding.
 7 Conclusion
-By analyzing Large Language Models (LLMs) as both psycholinguistic and neurolinguistic subjects, we find that LLMs excel at encoding linguistic form but show limitations in understanding meaning.
-Their grasp of meaning is largely dependent on statistical associations with form.
+This study adopts both psycho- and neuro-linguistic approaches to evaluating LLMs, revealing a distinction between linguistic performance and competence.
+Our results highlight the limitations of LLMs’ semantic understanding and the need for future research to move beyond statistical correlations toward more grounded language representations.
 By introducing a cognitive neuroscience perspective, along with semiotics, we hope will inspire further research to deepen our understanding of the language capabilities of LLMs.
 Limitations
-We didn’t include experiments covering more languages, which might limit the generalizability of our findings.
-The evaluation results for German are poor, which might be attributed to the issue of very long sentences in the DistilLingEval dataset.
-Due to limitations in computational resources, we did not perform our experiments on larger-size LLMs, which may lead our results to be biased towards small-scale models.
-Further studies should include larger models to validate and potentially generalize our findings.
+This study has several limitations that may impact the generalizability and comprehensiveness of our findings.
+First, we did not include experiments covering a wider range of languages, which restricts the cross-linguistic applicability of our results.
+Especially for the analysis and discussion in Section 5.3 on multilingual content, which highlights the necessity of constructing multilingual conceptual datasets.
+Second, the evaluation results for German are notably poor, potentially due to the presence of very long sentences in the DistilLingEval dataset, which might have introduced challenges for the models.
+This underscores the need for constructing syntactic minimal pair datasets for German.
+Additionally, our experiments were conducted using small-scale LLMs due to computational resource constraints.
+This may have introduced a bias in our findings, as larger-scale models could exhibit different behaviors.
+Future studies should explore larger models to validate and extend the generalizability of these results.
+Lastly, the COMPS dataset used for assessing conceptual understanding is not sufficiently fine-grained, as it is limited to only four types of conceptual relationships.
+A more granular dataset could provide deeper insights into the nuances of how LLMs encode and process meaning.
+Future work should address this limitation by incorporating more diverse and detailed datasets.
 Ethics Statement
-Our project may raise awareness within the computational linguistics community about those low-resource languages.
-Appendix A More Psycholinguistic and Neurolinguistic Paradigms’s Cognitive Science Background
+Our project has the potential to raise greater awareness within the computational linguistics community about the challenges faced by low-resource languages.
+By highlighting the unique linguistic features and limited computational tools available for these languages, we aim to inspire further research and the development of more inclusive language technologies that can better serve underrepresented linguistic communities.
+Appendix A Extended Cognitive Science Background
 Psycholinguistics often involves examining real-time language processing, linguistic knowledge storage, and language acquisition, using behavioral experimental methods such as reading times and eye-tracking.
@@
 It tries to extract information encoded in neural patterns (Kriegeskorte et al., 2006).
-It trains a classifier to predict properties of the stimulus (e.g.
+It trains a classifier to predict the properties of the stimulus (e.g.
 a particular image or word input), from the neural responses.
@@
 This dataset contains 49,340 sentence pairs, constructed using 521 concepts and 3,592 properties.
-Concepts in the pairs constitute 4 types or knowledge relationship: taxonomy, property norms, co-occurrence and random.
+Concepts in the pairs constitute 4 types of knowledge relationships: taxonomy, property norms, co-occurrence, and random.
 COMPS-ZH and COMPS-DE
@@
 This rigorous process was essential for preserving the integrity and reliability of the translated datasets.
-Appendix C More Detailed Minimal Pair Probing Results for English
-Figure 10: English minimal pair probing results.
-We averaged the results of 4 tasks from the same linguistic level, respectively, and show the result curves in Figure 10. Some noteworthy points from this figure include:
-1) The average performances on morphological tasks (purple curve) and on tasks from the semantics-syntax interface (yellow curve) are lower than those for syntactic tasks (blue curve), suggesting that more abstract linguistic levels are easier for LLMs to learn, which also aligns with He et al.
+Appendix C Model Details
+In our experiments, we used three open-source LLMs, two English-centric LLMs (Llama2 and Llama3), and one multilingual LLMs (Qwen) with a focus on English and Chinese.
+These models were trained on different amounts of English, Chinese, and German data (see Table 3).
+Resource Level Llama2 Llama3 Qwen
+English High High High
+Chinese Mid Mid High
+German Low Low Low
+Table 3: Resource level for different languages across three LLMs.
+Note the resource levels are qualitative assessments based on available information, as specific quantitative data is not provided by the developers.
+Llama2 and Llama3
+Llama2 (Touvron et al., 2023b) and Llama3 (AI, 2024) are two English-centric LLMs which represent an advanced iteration of the Llama foundation models developed by Meta AI (Touvron et al., 2023a).
+The Llama madels were trained on publicly available corpora predominantly in English.
+Despite this focus, Llama models are also exposed to a limited amount of multilingual data.
+Llama 1, for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages.
+Llama 2 expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data.
+Therefore, English-centric models harness multilingual abilities (Lai et al., 2023).
+In this work, we use Llama2-7B and Llama3-8B for our experiments.
+QWen
+QWen is a series of LLMs developed by Alibaba Inc.
+Bai et al.
 (2023).
-For the ellipsis task especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
+Qwen was trained on 2-3 trillion tokens of multilingual pretaining data.
+It is essentially a multilingual LLM with the focus on English and Chinese.
+We use the Qwen-7B model in our experiments.
+Appendix D Extended Neurolinguistic Results for English
+Some noteworthy points from Figure 5-(a) include:
+1) For the ellipsis task especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
 2) Among the conceptual tasks, the random relationship shows significantly higher accuracy compared to the other three conceptual relationships, suggesting that LLMs find it challenging to distinguish between similar concepts.
-Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much, but will learn concepts better and faster.
-Figure 11: Performance difference between Llama3 and Llama2.
-Figure 12: Layer-wise t-test results.
+Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much but will learn concepts better and faster.
+Figure 6: Top: Performance difference between Llama3 and Llama2.
+Bottom: Layer-wise t-test results.
 T-tests were first performed separately on each linguistic task, and then Stouffer’s Z-score method Stouffer et al.
@@
 This suggests that Llama3 comprehends sentences faster.
-Appendix D Multilingual Analysis
+Appendix E Extended Neurolinguistic Results for Multilingual Analysis
 1) Llama’s performance on Chinese.
@@
 This highlights differences in the resource allocation for the three languages.
+5) Form needs less data to capture compared to meaning.
+From Table 3, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
+Appendix F Extended Discussion
+Developmental difference between human and machine intelligence.
+From a perspective of developmental psychology, human kids typically acquire conceptual understanding before mastering grammar Bloom (2002); Tomasello (2005).
+Pinker (2009)’s semantic bootstrapping hypothesis posits that children initially learn vocabulary through semantic information and then use this semantic knowledge to infer syntactic structures.
+In contrast, our results indicate that LLMs learn grammar before meaning.
+Human intelligence is a combination of statistical inference and causal reasoning, whereas LLMs’ intelligence is more likely a result of statistical inference Tenenbaum et al.
+(2011); Gopnik and Wellman (2012); Lake et al.
+(2017).
+Given this nature, the fact that LLMs learn form first might be because grammatical patterns are easier to statistically capture compared to meaning.
+Symbol grounding problem and the quest for human-like intelligence
+In human language, the relationship between the signifier and the signified is often flexible and context-dependent, allowing for a more independent connection between syntax and semanticsHarnad (1990).
+Human cognitive development typically involves acquiring conceptual understanding first, followed by the learning of rules and syntax.
+In contrast, our study shows that LLMs grasp syntax before meaning, relying on statistical correlations within formal structures to infer semantic content.
+This difference highlights a fundamental divergence between human and machine intelligence, as LLMs do not possess an inherent understanding of meaning detached from the formal structures they analyze.
+These observations suggest that, for LLMs to develop human-like intelligence, they must transcend mere statistical pattern recognition.
+This will likely require the integration of world knowledge and grounded experiences that go beyond linguistic inputs.
+To achieve a more robust form of artificial intelligence that mirrors human cognition, models must be able to ground symbols in real-world contexts, establishing a basis for genuine understanding Tenenbaum et al.
+(2011); Lake et al.
+(2017) As it stands, the symbol grounding problem remains a significant barrier, and addressing it may be essential for constructing systems capable of true human-like reasoning and understanding.
 Figure 8: Detailed English decoding results on 6 models.
 Figure 9: Detailed Chinese decoding results on 6 models.
-Notice that the pink, orange, blue curves don’t denote morphology or semantics as those in English do.
+Notice that the pink, orange, and blue curves don’t denote morphology or semantics as those in English do.
 They are made just to make it easier to distinguish in the figure.
-All non-red curves represent grammatical tasks and red curves represent conceputal tasks.
+All non-red curves represent grammatical tasks and red curves represent conceptual tasks.
 Figure 10: Detailed German decoding results on 6 models. All non-red curves are grammatical tasks, and red curves are conceptual tasks.

### v3

@@
-Large Language Models as Neurolinguistic Subjects: Discrepancy in Performance and Competence for Form and Meaning
+Large Language Models as Neurolinguistic Subjects: Discrepancy between Performance and Competence
 Abstract
@@
 Traditional psycholinguistic evaluations often reflect statistical rules that may not accurately represent LLMs’ true linguistic competence.
-We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pair and diagnostic probing to analyze activation patterns across model layers.
+We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pairs and diagnostic probing to analyze activation patterns across model layers.
 This method allows for a detailed examination of how LLMs represent form and meaning, and whether these representations are consistent across languages.
@@
 Additionally, we introduce new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets.
+[^1] Code and data available here.
 1 Introduction
@@
 LLMs have drawn the attention of researchers from various fields, including linguistics, cognitive science, computer science, and neuroscience, who investigate how LLMs develop and exhibit these capabilities.
-Figure 1: Left: Research Question.
-Right: Psycholinguistic vs.
+Figure 1: Illustration of LLMs processing the same signified (meaning) across different signifiers (forms).
+Figure 2: Psycholinguistic vs.
 Neurolinguistic Paradigm.
-Both direct probability measurement and metalinguistic prompting can be considered psycholinguistic methods, while minimal pair probing He et al.
+Both direct probability measurement and metalinguistic prompting can be considered as psycholinguistic methods, while minimal pair probing He et al.
 (2024) and other diagnostic probing are neurolinguistic.
@@
 Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how form and meaning intertwist with each other.
-Is LLM’s understanding of language meaning merely a statistical outcome based on their grasp of language form?
+Is LLMs‘ understanding of language meaning merely a statistical outcome based on their grasp of language form?
 When different languages express a shared concept with distinct forms, do LLMs create similar representations for these variations?
@@
 Computing a model’s probability of generating two minimally different sentences is one way to address these concerns Hu and Levy (2023).
-The minimal difference between the two sentences (e.g.
-replacement of a single word) makes one sentence acceptable while the other is not Linzen et al.
+The minimal difference between the two sentences (e.g., replacement of a single word) makes one sentence acceptable while the other is not Linzen et al.
 (2016).
@@
 (2016); Tenney et al.
-(2019); Manning et al.
+(2019); Hewitt and Manning (2019); Manning et al.
 (2020).
@@
 A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$.
-Note that our focus is the performance of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, instead of the capability of the supervised classifier $g$.
-To eliminate the influence of $g$, we set up a “random” baseline, where we assign a random vector $r$ to each sentence to replace the hidden states $f_{l}(S)$ and calculate the performance of the classifier $g$ on the random vector $r$ as $\mathrm{Perf}(g,\mathcal{D}_{m})$, aiming to evaluate the inherent predictability of the probing classifier.
-Finally, the performance of the LLM $f$ is defined as follows:
-$\mathrm{Perf}(f,\mathcal{D}_{O})\triangleq\frac{\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})-\mathrm{Perf}(g,\mathcal{D}_{m})}{1-\mathrm{Perf}(g,\mathcal{D}_{m})}$ (1)
+Note that our focus is on evaluating the linguistic competence of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, rather than the capacity of the probing classifier $g$.
+As suggested by Hewitt and Liang (2019), even untrained or random representations can yield surprisingly high probing accuracy, raising concerns that the classifier may exploit dataset artifacts rather than meaningful representations.
+To control for the potential bias introduced by $g$, we construct a random embedding baseline.
+Specifically, for each sentence in the dataset, we assign a fixed random vector $r$, sampled from a Gaussian distribution with the same mean and standard deviation as the real model embeddings $f_{l}(S)$.
+Importantly, each sentence is consistently assigned the same random vector across occurrences, preserving instance-level identity that the probing classifier might exploit.
+This allows us to assess the extent to which task performance can be driven by superficial sentence-level cues rather than meaningful representations.
+We then compute $\mathrm{Perf}(g,\mathcal{D}_{m})$ by training $g$ on these random embeddings, which reflects the inherent predictability or “shortcut” potential of the probing task.
+Therefore, our performance score incorporates a correction factor based on this random baseline, defined as:
+$\mathrm{Perf}(f,\mathcal{D}_{O})\triangleq\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})\cdot(1+\frac{0.5-\mathrm{Perf}(g,\mathcal{D}_{m})}{0.5})$ (1)
+This formula applies a correction term, penalizing cases where the probing classifier performs well even on random embeddings.
+When $\mathrm{Perf}(g,\mathcal{D}_{m})=0.5$, the correction factor is 1; if the performance is higher, the factor shrinks toward 0, discouraging overfitting or trivial tasks; if it drops below 0.5, the factor exceeds 1, slightly amplifying the model’s score.
+This ensures that only meaningful representations in $f$ contribute to the final evaluation.
 4 Experiment Setup
@@
 Further dataset and model descriptions are in Appendix B and C.
+Figure 3: Psycholinguistic (meta and direct) and neurolinguistic performance across models and linguistic tasks.
+The x-axis represents different models and conditions (base and chat), while the y-axis categorizes linguistic tasks based on structural (syntax, morphology, syntax-semantics interface) and conceptual (meaning) levels.
 4.2 Setup for Psycholinguistic Analysis
@@
 Following Hu and Levy (2023), we use one prompt for a minimal pair to present both sentences at once.
-For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences and compare the probabilities assigned by the model to each answer option, “1” or “2”.
+For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences, and compare the probabilities assigned by the model to each answer option, “1” or “2”.
 For meaning tasks, we reformulate the property into a question and compare the probabilities of acceptable and unacceptable concepts as sentence continuations.
 Table 2 presents the prompts used in the experiments.
-Figure 2: Psycholinguistic (meta and direct) and neurolinguistic performance across models and linguistic tasks.
-The x-axis represents different models and conditions (base and chat), while the y-axis categorizes linguistic tasks based on structural (syntax, morphology, syntax-semantics interface) and conceptual (meaning) levels.
 4.3 Setup for Neurolinguistic Analysis
@@
 Probing Performance
-We use logistic regression as the probing classifier and F1 score as the evaluation metrics.
+We use logistic regression as the probing classifier and F1 score as the evaluation metric.
 The score for $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$ and $\mathrm{Perf}(g,\mathcal{D}_{m})$ is calculated as the average F1 score across 5 cross-validation folds.
@@
 5 Results
-Figure 3: (a) Averaged psycholinguistic (meta and direct) and neurolinguistic results across models and tasks.
-t-tests were conducted on the original (pre-averaging) results between base and chat models, with p-values annotated.
-(b) Psycholinguistic and neurolinguistic performance for form (morphology, syntax-semantics interface, and syntax) and meaning (concept).
-(c) Competence and performance gap drops after instruction tuning.
 5.1 Psycholinguistic vs Neurolinguistic
 Figure 3 shows the performance of LLMs across all linguistic tasks.
-Figure 3-(a) demonstrates the averaged performance of LLMs across models and 4 levels (syntax, morphology, syntax-semantics interfaces, concept).
-Figure 3-(b) presents the average performance of LLMs across form and meaning tasks for Direct, Meta, and Neuro [^1] We refer to minimal pair probing as Neuro for simplicity.
+Figure 4 demonstrates the averaged performance of LLMs across models and 4 levels (syntax, morphology, syntax-semantics interfaces, concept).
+Figure 5 presents the average performance of LLMs across form and meaning tasks for Direct, Meta, and Neuro [^2] We refer to minimal pair probing as Neuro for simplicity.
 methods.
@@
 Language performance and competence are distinct (Competence $>$ Performance).
-Figure 3 shows distinct results between language performance and competence.
+Figure 4 and 5 shows distinct results between language performance and competence.
 Moving from Meta → Direct → Neuro, the evaluation focus gradually shifts from language performance (task execution ability) to language competence (the underlying linguistic ability).
@@
 This indicates that when evaluating pure linguistic competence, LLMs perform well, but their performance drops when assessed in a task-based setting.
+Figure 4: Averaged psycholinguistic (meta and direct) and neurolinguistic results across models and tasks.
+t-tests were conducted on the original (pre-averaging) results between base and chat models, with p-values annotated.
 Tasks that emphasize language performance become more difficult, even if their language competence is high.
 For example, in the Neuro setting, performance on Syntax tasks reaches 97%, while in the Meta setting, it drops to 56.1%, showing a significant gap.
-This suggests that even when an LLM has strong competence in a given task, its performance can significantly decline when assessed under performance-oriented evaluation.
-Direct probability measurement might not be true competence assessment.
+This suggests that even when an LLM has strong competence in a given task, its performance can significantly decline when assessed under a performance-oriented evaluation.
+Figure 5: Psycholinguistic and neurolinguistic performance for form (morphology, syntax-semantics interface, and syntax) and meaning (concept).
+Direct probability measurement might not be a true competence assessment.
 As the Neuro method measures the internal representations of LLMs directly, it could serve as a reliable ground truth for estimating linguistic competence.
-Direct probability measurement falls short of achieving this ground truth in form assessment (especially for syntax and syntax-semantics-interface as shown in Figure 3-(b)).
+Direct probability measurement falls short of achieving this ground truth in form assessment (especially for syntax and syntax-semantics-interface as shown in Figure 5).
+Figure 6: Competence and performance gap drops after instruction tuning.
+Figure 7: t-SNE visualization of embedding differences between acceptable and unacceptable sentences, with red for syntax, purple for morphology, and yellow for the syntax-semantics interface.
 LLMs exhibit stronger mastery of form than meaning, regardless of performance or competence.
-As shown in Figure 3-(b), LLMs consistently perform better on form-related tasks than on meaning-related tasks.
+As shown in Figure 5, LLMs consistently perform better on form-related tasks than on meaning-related tasks.
 This trend holds regardless of whether the model is a base or chat version.
@@
 Neuro results between the base and chat versions of LLMs reveal that instruction fine-tuning does not significantly alter the language competence of the models (t-test between Neuro-Base vs.
-Neuro-Chat as shown in Figure 3-(a)).
+Neuro-Chat as shown in Figure 4).
 With instruction fine-tuning (chat versions of LLMs), Meta performance on form improves significantly while meaning understanding remains stable.
-Figure 3-(c) illustrates that after instruction tuning, the competence-performance gap (Neuro-Meta) significantly decreases for form-related tasks, while the change for meaning-related tasks remains relatively small.
+Figure 6 illustrates that after instruction tuning, the competence-performance gap (Neuro-Meta) significantly decreases for form-related tasks, while the change for meaning-related tasks remains relatively small.
 This indicates that fine-tuning with well-designed instructions helps LLMs improve their performance on form-related tasks, bringing them closer to their underlying competence.
@@
 This indicates that more optimized information access strategies can enhance the external performance of language models, particularly for form-related tasks.
-5.2 Neurolinguistic Analysis [^2] Raw results for English, Chinese, and German can be found in Figure 8, 9 and 10 in the Appendix.
-Figure 4: t-SNE visualization of embedding differences between acceptable and unacceptable sentences, with red for syntax, purple for morphology, and yellow for the syntax-semantics interface.
+Figure 8: (a) Neurolinguistic probing performance for 16 tasks in Llama-2, including 4 syntax tasks, 4 morphology tasks, 4 syntax-semantics interface tasks, and 4 conceptual tasks.
+(b) Average probing performance across the four linguistic categories in Llama-2.
+(c) Mean performance comparison between form-related tasks (syntax, morphology, syntax-semantics interface) and meaning-related tasks (concept), aggregated across all six models.
+Figure 9: Feature learning saturation layer (defined as the first layer reaching 95% of peak performance) and the layer of maximum performance.
+Figure 10: Difference in probing performance between base and instruction-tuned models across all layers.
+5.2 Neurolinguistic Analysis [^3] Raw results for English, Chinese, and German can be found in Figure 15, 16 and 17 in the Appendix.
 Layer-wise unsupervised dynamics reveal gradual emergence of form features
@@
 This pattern is consistent across all six models, suggesting a universal characteristic of LLMs.
-Moreover, as illustrated in Figure 5-(a),(b), the encoding performance progressively declines from more structural tasks to more semantic tasks—spanning syntax, morphology, the syntax-semantic interface, and finally conceptual understanding.
+Moreover, as illustrated in Figure 8-(a),(b), the encoding performance progressively declines from more structural tasks to more semantic tasks, spanning syntax, morphology, the syntax-semantic interface, and finally conceptual understanding.
 This highlights that LLMs encode features less effectively as the tasks shift from structure-focused to meaning-focused.
-Figure 5: (a) Neuro probing results for 12 tasks in Llama-2 (4 syntax, 4 morphology, 4 syn.-sem.
-inter., 4 concept).
-(b) Averaged performance across four linguistic levels in Llama-2.
-(c) Averaged performance on form vs.
-meaning across all six models.
-(d) Feature learning saturation (95% of peak) and maximum layer.
-(e) Perf.
-difference between base and instruction-tuned LLMs.
-(f) Neuro probing results for Chinese and German.
-(g) Correlation between meaning competence and form competence.
 LLMs encode form earlier than meaning.
 We compute the feature learning saturation and maximum layers for all 12 grammatical tasks and 4 conceptual tasks, averaging them to represent form and meaning, respectively.
-As shown in Figure 5-(d), the saturation and maximum layers for meaning are generally higher than those for form across all six models.
+As shown in Figure 9, the saturation and maximum layers for meaning are generally higher than those for form across all six models.
 This suggests that LLMs stabilize their encoding of grammatical features before conceptual features.
 Instruction tuning has minimal impact on the internal linguistic representations.
-As Figure 5-(e) shows, performance differences for form and meaning remain near zero across all layers, indicating that instruction tuning minimally impacts internal linguistic representations, consistent with our psycholinguistic vs.
+As Figure 10 shows, performance differences for form and meaning remain near zero across all layers, indicating that instruction tuning minimally impacts internal linguistic representations, consistent with our psycholinguistic vs.
 neurolinguistic analysis.
 5.3 Multilingual analysis
+Figure 11: Neuro probing results for English, Chinese, and German.
 How does LLMs’ understanding of meaning change when the form (language) varies?
@@
 Our previous results suggest that instruction tuning has little influence on the internal representations.
-Therefore we focus on the base LLMs here.
-Cross-lingual Pattern
-From Figure 5-(f), for all models and languages, form consistently achieves higher performance than meaning, indicating it’s easier for LLMs to make a stronger grasp of structural elements compared to conceptual comprehension.
-Extended multilingual results can be found in Appendix E.
-Meaning competence is correlated to form competence.
-Figure 5-(g) illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs.
-The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
+Therefore, we focus on the base LLMs here.
+From Figure 11, for all models and languages, form consistently achieves higher performance than meaning, indicating it’s easier for LLMs to make a stronger grasp of structural elements compared to conceptual comprehension.
+Extended multilingual analysis can be found in Appendix E.
 6 Discussion
-Language performance vs. competence: using neurolinguistic probing to reveal deeper competencies.
-Our results demonstrate distinct patterns between the neurolinguistic and psycholinguistic evaluation methods.
-We observed that LLM performance evaluated using the Meta method was significantly lower than that of Direct, while Direct also performed notably worse than the Neuro method.
-This finding extends Hu and Levy (2023), who compared direct probability measurement (Direct) and meta-linguistic prompting (Meta) and argued that performance does not always reflect competence.
-While Hu and Levy (2023) suggested that Direct probability measurement is a more reliable approach for assessing competence, our findings reveal that there is still a significant gap between Direct and Neuro probing, which directly examines internal representations.
-Hence, we recommend Neuro probing for evaluating competence.
-Neuro probing further validates the discrepancy between performance and competence.
-While Meta results underperform, this does not necessarily indicate that the LLMs lack the underlying linguistic competence.
+Language performance vs. competence: probing reveals deeper linguistic understanding than direct probability.
+Our results demonstrate that neurolinguistic probing uncovers linguistic competencies in LLMs that are not captured by psycholinguistic methods.
+While Meta performs the worst and Direct performs better, Neuro consistently outperforms both, revealing a systematic underestimation of competence when relying on output-based evaluations.
+Hu and Levy (2023) argued that Direct probability measurement, being more intrinsic than metalinguistic prompting, better reflects competence.
+However, our findings show that even Direct falls short of revealing the full extent of LLMs’ linguistic capabilities.
+Direct relies on the final output layer, which is highly optimized for next-word prediction and thus entangled with task-specific objectives.
+Prior studies Hewitt and Manning (2019); Liu et al.
+(2019) have shown that syntactic and general linguistic information is often better represented in intermediate layers than in the final layer.
+Waldis et al.
+(2024) also emphasized that output correctness is an insufficient indicator of linguistic understanding, advocating for probing internal representations.
+Our t-SNE visualizations corroborate this: clear linguistic clusters emerge in intermediate layers but dissolve in the final layer, reinforcing the view that the last layer is not optimal for assessing competence.
+These findings suggest that Direct, while more grounded than prompting, is still a limited proxy for internal knowledge.
+In contrast, neurolinguistic probing inspects internal activation patterns across layers and tasks, uncovering the underlying representational structure of form and meaning, and further validates the discrepancy between performance and competence.
+On the other hand, while Meta results underperform, this does not necessarily indicate that the LLMs lack the underlying linguistic competence.
 Instead, it may reflect limitations in information access, as suboptimal prompts can prevent models from exhibiting their full capabilities.
 Specifically, prompting failures do not always equate to a lack of encoded knowledge.
-This observation is consistent with previous studies Firestone (2020); Lampinen (2024) that emphasize the importance of carefully controlling performance conditions for both humans and machines.
-methods can often reveal deeper competencies within the models.
+This aligns with prior work Firestone (2020); Lampinen (2024) emphasizing the need to distinguish performance conditions from underlying ability.
+Thus, we argue that probing, particularly when applied layer-wise, provides a more accurate and comprehensive assessment of linguistic competence than Direct probability alone.
 Form and meaning: observations from Saussure’s semiotics
@@
 This suggests that LLMs’ understanding of meaning might rely heavily on form, with conceptual representation anchored to formal structures rather than independent meaning comprehension.
-These results offer a semiotic and neurolinguistic explanation for LLMs’ long-standing issue of generating “confidently incorrect” responses, i.e.
-hallucinations Ji et al.
+These results offer a semiotic and neurolinguistic explanation for LLMs’ long-standing issue of generating “confidently incorrect” responses, i.e., hallucinations Ji et al.
 (2023).
@@
 By highlighting the unique linguistic features and limited computational tools available for these languages, we aim to inspire further research and the development of more inclusive language technologies that can better serve underrepresented linguistic communities.
-Appendix A Extended Cognitive Science Background
+Acknowledgements
+We thank the anonymous reviewers for their valuable advice and feedback.
+This research was supported by DFG (German Research Foundation) grant SCHU 2246/14-1 and Munich Center for Machine Learning (MCML).
+Appendix A Neuro Probing’s Cognitive Science Background
 Psycholinguistics often involves examining real-time language processing, linguistic knowledge storage, and language acquisition, using behavioral experimental methods such as reading times and eye-tracking.
@@
 They use electrocorticography (ECoG) to record neural responses from subjects who are listening to speech.
-Leveraging decoding analysis, they were able to differentiate between neural patterns induced by attended speech and those elicited by background speech, (to be ignored by the test persons), thereby highlighting the perceptual differences between the two speech stimuli conditions.
+Leveraging decoding analysis, they were able to differentiate between neural patterns induced by attended speech and those elicited by background speech (to be ignored by the test persons), thereby highlighting the perceptual differences between the two speech stimulus conditions.
 While psycholinguistic approaches provide valuable insights into LLMs’ functional capabilities, they often fall short in revealing the underlying mechanisms of language processing.
@@
 CLiMP (Xiang et al., 2021) is a corpus of Chinese grammatical minimal pairs consisting of 16 datasets, each containing 1,000 sentence pairs.
-CLiMP covers 9 major Chinese language phenomena in total, less than the BLiMP dataset due to the less inflectional nature of Mandarin Chinese.
+CLiMP covers 9 major Chinese language phenomena in total, fewer than the BLiMP dataset due to the less inflectional nature of Mandarin Chinese.
 The vocabulary of the CLiMP dataset is based on the translation of the BLiMP dataset, with words and features specific to Chinese added.
@@
 In summary, out of 521 concepts, manual corrections were made to 57 entries in the Chinese dataset and 49 in the German dataset.
-Similarly, from 3,592 properties, 713 required manual corrections in the Chinese dataset, and 512 in the German dataset.
+Similarly, out of 3,592 properties, 713 required manual corrections in the Chinese dataset, and 512 in the German dataset.
 This rigorous process was essential for preserving the integrity and reliability of the translated datasets.
 Appendix C Model Details
-In our experiments, we used three open-source LLMs, two English-centric LLMs (Llama2 and Llama3), and one multilingual LLMs (Qwen) with a focus on English and Chinese.
+In our experiments, we used three open-source LLMs, two English-centric LLMs (Llama2 and Llama3), and one multilingual LLM (Qwen) with a focus on English and Chinese.
 These models were trained on different amounts of English, Chinese, and German data (see Table 3).
@@
 Llama 1, for example, is pretrained on an extensive scale of corpora comprising over 1.4 trillion tokens, of which less than 4.5% constitute multilingual data from 20 different languages.
-Llama 2 expands this linguistic diversity, featuring 27 languages each representing more than 0.005% of the pertaining data.
+Llama 2 expands this linguistic diversity, featuring 27 languages, each representing more than 0.005% of the pertaining data.
 Therefore, English-centric models harness multilingual abilities (Lai et al., 2023).
@@
 (2023).
-Qwen was trained on 2-3 trillion tokens of multilingual pretaining data.
-It is essentially a multilingual LLM with the focus on English and Chinese.
+Qwen was trained on 2-3 trillion tokens of multilingual pre-training data.
+It is essentially a multilingual LLM with a focus on English and Chinese.
 We use the Qwen-7B model in our experiments.
-Appendix D Extended Neurolinguistic Results for English
+Appendix D Supplemented Results for English
 Some noteworthy points from Figure 8-(a) include:
-1) For the ellipsis task especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
+1) For the ellipsis task, especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
 2) Among the conceptual tasks, the random relationship shows significantly higher accuracy compared to the other three conceptual relationships, suggesting that LLMs find it challenging to distinguish between similar concepts.
-Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much but will learn concepts better and faster.
+Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much, but will learn concepts better and faster.
 Figure 12: Top: Performance difference between Llama3 and Llama2.
@@
 The t-test statistics in Figure 12 further support these results.
-Figure 7: T-test results between Llama2 and Llama3 feature learning saturation layer. The symbols ‘***’, and ‘*’ denote t-test p-values less than 0.001, and 0.05, respectively.
+Figure 13: T-test results between Llama2 and Llama3 feature learning saturation layer. The symbols ‘***’, and ‘*’ denote t-test p-values less than 0.001 and 0.05, respectively.
 Figure 13 compares the feature learning saturation layers between Llama2 and Llama3.
-The results for form learning (blue bars) do not differ significantly between Llama2 and Llama3 and weakly significantly between Llama2_chat and Llama3_chat.
+The results for form learning (blue bars) do not differ significantly between Llama2 and Llama3, and weakly significantly between Llama2_chat and Llama3_chat.
 However, the results for meaning learning (blue bars) are both highly significant, indicating that Llama3 requires fewer layers to encode conceptual features than Llama2.
 This suggests that Llama3 comprehends sentences faster.
-Appendix E Extended Neurolinguistic Results for Multilingual Analysis
-1) Llama’s performance on Chinese.
+Appendix E Supplemented Results for Multilingual Analysis
+Figure 14: Correlation between meaning competence and form competence.
+Meaning competence is correlated to form competence.
+Figure 14 illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs.
+The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
+Llama’s performance on Chinese.
 Despite Chinese not being the primary training language of the Llama2 models, they still perform well in encoding Chinese form/grammar.
 However, both Qwen, which is primarily trained on Chinese, and the Llama models show relatively poor performance in understanding Chinese meaning/concepts.
-2) Improvement in Llama3 for Chinese Semantics.
-Llama3 shows some improvement over Llama2 in understanding Chinese semantics, as indicated by the slightly higher red curve in the middle row.
-The improvement in syntactic understanding is minimal.
-3) Qwen’s Faster Syntax Learning but Slower Semantic Encoding for Chinese.
+Improvement in Llama3 for Chinese Semantics.
+Llama3 shows some improvement over Llama2 in understanding Chinese semantics, as indicated by the slightly higher red curve in the middle row. The improvement in syntactic understanding is minimal.
+Qwen’s Faster Syntax Learning but Slower Semantic Encoding for Chinese.
 Compared to the Llama models, Qwen learns Chinese grammar faster, as indicated by the sharper rise of the blue curve.
 However, it encodes Chinese semantics more slowly, evidenced by the larger gap between the form and meaning curves in the early layers.
-4) Poor Performance for German.
+Poor Performance for German.
 For German, a low-resource language, all three models perform poorly.
@@
 This highlights differences in the resource allocation for the three languages.
-5) Form needs less data to capture compared to meaning.
+Form needs less data to capture compared to meaning.
 From Table 3, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
-Appendix F Extended Discussion
+Appendix F Supplemented Discussion
 Developmental difference between human and machine intelligence.